# Learning Agility Adaptation for Flight in Clutter

### v1

@@ IV-C System Overview
At the uppermost end of the system, the vehicle obtains partial observations of the environment via external sensors, which are used to provide state feedback (e.g., via the visual-inertial odometry (VIO)) and to construct local occupancy maps (e.g., from the depth camera). The local map and state estimates are fed to the RL policy and trajectory planner. We choose to use the occupancy map as the input of the RL policy instead of the raw sensory data because the training is performed in simulation, and such an approach deals with sensor noise through a model-based approach, i.e., the employed mapping algorithm, thus bridging the gap between simulation and reality in RL training.
At each time step $t$, the agility policy outputs a high-level planning instruction $v^{\dagger}_{t}$. The trajectory planner generates a trajectory $u$ in a horizon $[t,t+T]$ conditioned on $v^{\dagger}_{t}$. The discretization of the trajectory is treated as the action $a_{t}$ of the POMDP. The trajectory is tracked by a low-level controller and executed by the vehicle’s inner-loop controller. In our definition, the unknown environment as well as low-level controllers are treated as dynamics of the POMDP.
@@
V Reinforcement Learning for Agility Policy
In this section, we detail how to efficiently and effectively learn the agility policy. Eq. (5) implies an approach to learn the agility policy, which treats the trajectory planner as part of the environment dynamics. An auxiliary POMDP is considered to learn $\pi^{\dagger}$, in which the action space is $V^{\dagger}$. In the following subsections, we present the implementation of each part of RL, where a couple of pre-training-fine-tuning reward functions are the keys to a successful training.
@@ V-A Policy Representation
We represent policy and value functions using a branch of partially shared neural networks, as illustrated in Fig. 2. The networks share a perception encoder consists of convolution neural networks (CNNs) and a fusion layer implemented with multilayer perceptron (MLP). We implement the CNN encoder with 4 layers whose [channel number, kernel size, stride, padding] are [5,5,3,2], [32,3,2,1], [48,3,2,1], and [64,3,1,0]. The result of the CNN encoder is flattened and concatenated into the inputs of the MLP fusion layer. The output dimension of the MLP layer is 64. The actor and critic networks are both implemented as MLPs with 2 hidden layers having 64 and 32 units, respectively.
@@ V-B Observation Space
As shown by (5), in principle, $\pi^{\dagger}$ should be conditioned on the observations and the goal of navigation. However, as illustrated in Fig. 2, we employ a more practical input which consists of the local 3D occupancy map $\boldsymbol{m}_{t}$, the current velocity $\boldsymbol{v}_{t}$, the current tracking error $\boldsymbol{e}_{t}\coloneqq\boldsymbol{p}_{t}-\hat{\boldsymbol{p}}_{t}$, where $\boldsymbol{p}_{t}$ and $\hat{\boldsymbol{p}}_{t}$ are the true and desired position of the vehicle, respectively, the agility decision at the last time step ${v}^{\dagger}_{t-1}$, and a pre-planned trajectory at the current time step generated according to ${v}^{\dagger}_{t-1}$ and $o_{t}$, which will not be executed by the controller.
The pre-planned trajectory serves two purposes. First, it expresses the current position of the vehicle in the local map as well as the local target chosen by the trajectory planner according to the global goal $g$. Second, since agility has little effect on the spatial distribution of the trajectory, we can use this pre-planned trajectory as a spatial approximation of the trajectory to be planned, and therefore alert the perceptual part of the policy as to which regions in the local map are important.
We correlate the trajectory and the 3D local map by ’drawing’ the former on the latter, as illustrated by Fig. 2. Specifically, each cell in the occupancy map is assigned one of four states: free, occupied, unknown, and on_trajectory. The state unknown is set to trigger perception-aware behaviors. For more efficient and easier learning, we sample the trajectory with time interval $\delta t$, keeping only the x-y 2D cut of the 3D map corresponding to the sampled point on the trajectory as input to the network, which is illustrated in Fig. 2.
V-C Termination
We define two cases of termination. One is when the vehicle is in an unsafe state, i.e., the vehicle is judged in the emergency_stop or collided state, which is determined by the state machine in a modern motion planning system [7, 17]. The other occurs when the planning solution procedure does not finish in a valid time, which is often due to too high level of agility allowed in the previous time steps, so that the vehicle is too close to a suddenly appearing obstacle under the perception latency and thus the planner fails to plan a feasible trajectory.
Figure 3: Example illustration of the naive estimate of the danger level. The blue blocks are the perceived local occupancy map. (a) A safe case where small number of obstacles are far from the vehicle. (b) A mild case. (c) A dangerous case where an obstacle is close to the vehicle.
@@ V-D Reward Function
As described in section III, we design the reward function to maximize agility and minimize collision loss. For practical considerations, we further divide the reward function into four parts: the agility term, the smoothing term, the tracking error term, and the (sparse) collision term:
$r=r_{\textnormal{agility}}+r_{\textnormal{smoothing}}+r_{\textnormal{error}}+r_{\textnormal{danger}},\vspace{-0.1cm}$ (6)
@@
$r_{\textnormal{smoothing}}=-\lambda_{\textnormal{smoothing}}\|v_{t}^{\dagger}-v_{t-1}^{\dagger}\|^{2},\vspace{-0.2cm}$ (7)
$r_{\textnormal{error}}=-\lambda_{\textnormal{error}}\min\left\{\|\boldsymbol{e}_{t}\|,e_{\max}\right\}^{2},\vspace{-0.1cm}$ (8)
$r_{\textnormal{danger}}=\begin{cases}-\lambda_{\textnormal{danger}}\|\boldsymbol{v}_{t}\|^{2}&\textnormal{(when the episode is terminal)}\\ 0&\textnormal{(otherwise)}\\ \end{cases},\vspace{-0.1cm}$ (9)
@@
However, since $r_{\textnormal{danger}}$ is highly sparse and stochastic, this form of reward is difficult to learn with for general RL algorithms, especially when embedded in a complex modern trajectory planner. Our key finding here is that with a two-stage reward design in a pre-training-fine-tuning training pipeline, learning can take place efficiently and effectively. The only term that differs between the first and second stages is $r_{\textnormal{agility}}$, which is detailed in the following.
@@ V-D1 A human knowledge-based, dense reward function for pre-training
The agility reward in the first stage is defined as follows:
$r_{\textnormal{agility}}=\begin{cases}\lambda_{\textnormal{agility}}^{1}(\phi_{t}^{1}-\|\boldsymbol{v}_{t}\|)&\left(\phi_{t}^{2}>2\right)\\ \lambda_{\textnormal{agility}}^{2}(\|\boldsymbol{v}_{t}\|-\phi_{t}^{1})&\left(\phi_{t}^{2}<1\right)\\ \lambda_{\textnormal{agility}}^{3}\|\boldsymbol{v}_{t}\|&\textnormal{(otherwise)}\\ \end{cases},\vspace{-0.1cm}$ (10)
where $\lambda_{\textnormal{agility}}^{i}>0$ for $i\in\left\{1,2,3\right\}$, $\lambda_{\textnormal{agility}}^{2}>\lambda_{\textnormal{agility}}^{3}$, and $\phi_{t}^{i}$s, for $i\in\left\{1,2\right\}$, are handcrafted feature values encoding the obstacle distribution. Specially, $\phi_{t}^{i}$s are normalized linear combinations of the distance to the nearest obstacle, volume of the obstacle, and number of obstacles at current time, some example of which are illustrated in Fig. 3.
In (10), $\phi_{t}^{1}$ is designed to shape the reward function smoother with a zero mean, while $\phi_{t}^{2}$ is a naive estimate of the level of danger by human knowledge. The first case in (10) serves as a dense smoothing of the collision penalty.
@@ V-D2 An objective reward for fine-tuning
In the second stage, we restore the intent of the agility reward, which is to encourage greater agility, as follows:
$r_{\textnormal{agility}}=\lambda_{\textnormal{agility}}^{3}\|\boldsymbol{v}_{t}\|.\vspace{-0.1cm}$ (11)
@@ VI-A System Setup
We choose a state-of-the-art model-based trajectory planner, EGO-planner [17], as the backbone. The agility constraint is imposed in the form of $\left\|\boldsymbol{v}\right\|\leq\bar{v}$. The local mapping algorithm is implemented based on [29]. The vehicle is equipped with a forward-facing camera that can measure depth with a FOV of ${}^{\circ}$ (horizontal) $\times$ ${}^{\circ}$ (vertical) and a valid/trusted depth range from $0.28$m to $5$m. The refreshing frequency of the depth data is 15 Hz. A PID controller running at 50 Hz is used for position tracking.
The agility policy runs at 10 Hz. Conceptually, trajectory planning is required after each agility policy output, but too frequent replanning increases the situations of triggering the emergency_stop state in practice, as well as increasing the computational overhead. Therefore, we design a criterion to determine whether replanning is imposed by the agility policy output, and if this criterion is not met, the planner follows its original implementation of replanning rules. In particular, if $v_{t}^{\dagger}-v_{t-1}^{\dagger}\in\left[-0.3,0.5\right]$m/s, we do not impose a replanning of the trajectory planner.
@@ VI-B Training Environment
We train the policy in a customized simulator. Unlike general simulation environments, the simulator does not stop the clock while solving for the action (i.e., trajectory planning) in our simulation environment. Meanwhile, the sensory data is updated asynchronously with the action. This is to simulate one of the most important factors limiting agility, the perception latency, i.e., the time interval between perception and execution of an action. Another crucial factor that limits agility, the imperfect performance of trajectory tracking, is also embedded in the environment by setting up the actual PID tracking controller for the vehicle, as described earlier.
@@ VI-C Training Implementation
Figure 5: Velocity distribution along the trajectory in different setups. The dark green curves are the reference trajectory. The colorful curves are the trajectory that the vehicle passes over, where the color represents the velocity. (a) An example result of the proposed approach. (b) An example result of the constant agility ${v}^{\dagger}=2$m/s. (c) An example result of the intermediate model before fine-tuning.
Figure 6: Statistical results of different setups. Success rates are computed on 50 trials for each setup. The light blue dashed lines connect the statistics at different levels of constant agility, indicating the inherent capability of the system. Average velocities are computed as the average of success trials in the 50 trials.
@@ VII-A Simulation Experiments
We benchmark three approaches in the simulation experiments: (i) the proposed ’agility policy + model-based planner backbone’, (ii) the backbone planner EGO-planner with constant agility (i.e., constant ${v}^{\dagger}$), and (iii) the handcrafted cost function-based (non-learning) environmental adaptive planner EVA-planner [23], where the default parameters in the cost function are used, and all parameters that EGO-planner shares with it are set equal to those of EGO-planner.
@@
VII-A1 Comparison of the proposed framework with constant agility baselines and ablation tests
The first two setups are evaluated in a challenging environment as shown in Fig. 5, where obstacles of different densities, sizes, and appearances are unevenly distributed. The evaluation environment is never seen when training. In Fig. 6, we report the statistical results of traversal velocity in success cases and the success rate which is defined as the proportion of trials that complete the entire track. A trial is judged to be ’terminal’ with the same criteria as during training. In Fig. 5, we visualize the trajectories and velocities at every moment of the proposed approach and the baselines. Since the trajectory generation backbone of EVA-planner cannot always effectively generate feasible trajectory in such a complex environment, it is not fair to include it in this statistical result, so we ignore the results of EVA-planner in Fig. 6.
The statistics show that as ${v}^{\dagger}$ is set higher, the success rate decreases, creating a sloping downward performance curve in Fig. 6. Such performance curves capture the inherent performance of the integration of physical (e.g., sensors equipped) and algorithmic (e.g., the planning and control framework) systems. Although under the performance constraints of the system, by dynamically configuring the agility based on observations, the vehicle combines the exploitation of agility and safety guarantees. Specifically, it can be seen in Fig. 6 that the policy learned with both the reward stages achieves similar success rate as the lowest agility case where ${v}^{\dagger}=1.5$m/s, but also exhibits considerable traversal speeds to efficiently complete the track. In contrast, a relatively aggressive constant agility ${v}^{\dagger}=2$m/s can frequently lead to emergency stops or collisions in areas with dense obstacles or local occlusion, as shown in Fig. 5.
In Fig. 6, we also show the performance of the policies trained only with either the first stage (reward (10)) or the second stage (reward (11)). The results indicate that although the final policy can successfully improve the overall performance of the system, the policy trained with only the first stage does not exhibit substantial improvement. This result is not so surprising since the reward scheme of the first stage of training does not reflect the objective values and impose inaccurate human knowledge. On the other hand, the necessity of the first training stage is reflected by the inferiority of the performance rendered by the policy trained directly with sparse (but objective) rewards in Fig. 6.

### v2

@@ IV-C System Overview
At the uppermost end of the system, the vehicle obtains partial observations of the environment via external sensors, which are used to provide state feedback (e.g., via the visual-inertial odometry (VIO)) and to construct local occupancy maps (e.g., from the depth camera). The state estimates and local map are fed to the policy $\pi^{\dagger}$ and trajectory planner. We choose to use the occupancy map as the input of the RL policy instead of the raw sensory data because the training is performed in simulation, and such an approach deals with sensor noise through a model-based approach, i.e., the employed mapping algorithm, thus bridging the gap between simulation and reality in RL training.
At each time step $t$, the policy $\pi^{\dagger}$ outputs a high-level planning instruction $v^{\dagger}_{t}$. The trajectory planner generates a trajectory $u$ in a horizon $[t,t+T]$ conditioned on $v^{\dagger}_{t}$, which is treated as the action $a_{t}$ of the POMDP. The trajectory is tracked by a low-level controller and executed by the vehicle’s inner-loop controller. In our definition, the unknown environment as well as tracking and actuating controllers are treated as dynamics of the POMDP.
@@
V Reinforcement Learning for the Outer-loop Policy
In this section, we detail how to effectively learn the outer-loop policy. Eq. (5) implies an approach to learn this policy, which treats also the trajectory planner as part of the environment dynamics. An auxiliary POMDP is considered to learn $\pi^{\dagger}$, in which the action space is $V^{\dagger}$. In the following subsections, we present the implementation of each part of the RL policy.
@@ V-A Policy Representation
We represent policy and value functions using a branch of partially shared neural networks, as illustrated in Fig. 2. The networks share a perception encoder consists of convolution neural networks (CNNs) and a fusion layer implemented with multilayer perceptron (MLP). We implement the CNN encoder with 4 layers whose [channel number, kernel size, stride, padding] are [5,5,3,2], [32,3,2,1], [48,3,2,1] and [64,3,1,0]. The result of the CNN encoder is flattened and concatenated into the inputs of the MLP fusion layer. The output dimension of the MLP layer is 64. The actor and critic networks are both implemented as MLPs with 2 hidden layers having 64 and 32 units, respectively.
@@ V-B Observation Space
As implied by (5), in principle, $\pi^{\dagger}$ should be conditioned on the observations and the goal of navigation. However, for more effective learning, we design a specialized input space. As illustrated in Fig. 2, the input consists of the 3D local occupancy map $\boldsymbol{m}_{t}$, the current velocity $\boldsymbol{v}_{t}$, the current tracking error $\boldsymbol{e}_{t}\coloneqq\boldsymbol{p}_{t}-\hat{\boldsymbol{p}}_{t}$, where $\boldsymbol{p}_{t}$ and $\hat{\boldsymbol{p}}_{t}$ are the true and desired position of the vehicle, respectively, the decision at the last time step ${v}^{\dagger}_{t-1}$, and a pre-planned trajectory at the current time step generated according to ${v}^{\dagger}_{t-1}$ and $o_{t}$, which will not be executed by the controller.
The pre-planned trajectory serves two purposes. First, it expresses the current position of the vehicle in the local map as well as the local target chosen by the trajectory planner according to the global goal $g$. Second, since the speed component of the trajectory does not have a significant effect on the spatial distribution of the trajectory, we can use this pre-planned trajectory as a spatial approximation of the trajectory to be planned, and therefore alert the perceptual part of the policy as to which regions in the local map are important.
We correlate the trajectory and the 3D local map by ’drawing’ the former on the latter, as illustrated by Fig. 2. Specifically, each cell in the occupancy map is assigned one of four states: free, occupied, unknown, and on_trajectory. The state unknown is set to trigger perception-aware behaviors. For more efficient and easier learning, we sample the trajectory with time interval $\delta t$, keeping only the x-y 2D cuts of the 3D map corresponding to the sampled points on the trajectory as input to the network, which is illustrated in Fig. 2.
V-C Early Termination
We define two cases of termination. One is when the vehicle is in an unsafe state, i.e., the vehicle is judged in the emergency_stop or collided state, which is determined by the state machine in a modern motion planning system [6, 12]. The other occurs when the planning solution procedure does not finish in a valid time, which is often due to too high level of aggressiveness allowed in the previous time steps, so that the vehicle is too close to a suddenly appearing obstacle under the perception latency and thus the planner fails to plan a feasible trajectory.
@@ V-D Reward Function
As described in section III, we design the reward function to maximize traversal speed and minimize collision loss. For practical considerations, we further divide the reward function into four parts: the speed term, as follows:
$r=r_{\textnormal{speed}}+r_{\textnormal{smoothing}}+r_{\textnormal{error}}+r_{\textnormal{danger}},$ (6)
@@
$r_{\textnormal{smoothing}}=-\lambda_{\textnormal{smoothing}}\|v_{t}^{\dagger}-v_{t-1}^{\dagger}\|^{2},$ (7)
$r_{\textnormal{error}}=-\lambda_{\textnormal{error}}\min\left\{\|\boldsymbol{e}_{t}\|,e_{\max}\right\}^{2},$ (8)
$r_{\textnormal{danger}}=\begin{cases}-\lambda_{\textnormal{danger}}\|\boldsymbol{v}_{t}\|^{2}&\textnormal{(when the episode is terminal)}\\ 0&\textnormal{(otherwise)}\\ \end{cases},$ (9)
@@
However, $r_{\textnormal{danger}}$ is highly sparse and stochastic, where the stochasticity is derived from the complex modern trajectory planner. This distribution of reward can cause difficulties when learning the policy with general RL algorithms. One way to relax the early-termination penalty and make it denser is to add the distance to obstacles into the reward function. However, although this may be effective for an end-to-end policy that directly outputs a low-level command [13], this kind of geometric metric cannot evaluate the risk of early termination in our framework, as the spatial distribution of the vehicle is mainly determined by the planned trajectory where smaller distances to obstacles do not necessarily imply that collisions are more likely to occur [12]. In principle, we should only settle for penalizing as an episode does terminate, which is causally relevant to some implicit factors such as perception latency and tracking failure that make overly aggressive trajectories prone to cause collisions.
Our key finding here is that with a two-stage reward, learning can take place efficiently and effectively. The only term that differs between the first and second stages is $r_{\textnormal{speed}}$, which is detailed in the following.
@@ V-D1 A human knowledge-based, dense reward function for pre-training
The speed reward in the first stage, where $v_{t}^{\dagger}$ is imposed to be larger or equal to a certain constant value when training (e.g., 1m/s), is defined as follows:
$r_{\textnormal{speed}}=\begin{cases}\lambda_{\textnormal{speed}}^{1}(\phi_{t}^{1}-\|\boldsymbol{v}_{t}\|)&(\phi_{t}^{2}>\lambda_{\phi}^{1})\\ \lambda_{\textnormal{speed}}^{2}(\|\boldsymbol{v}_{t}\|-\phi_{t}^{1})&(\phi_{t}^{2}<\lambda_{\phi}^{2})\\ \lambda_{\textnormal{speed}}^{3}\|\boldsymbol{v}_{t}\|&\textnormal{(otherwise)}\\ \end{cases},$ (10)
where $\lambda_{\textnormal{speed}}^{i}>0$ for $i\in\left\{1,2,3\right\}$ and $\lambda_{\phi}^{i}$ for $i\in\left\{1,2\right\}$ are hyperparameters such that $\lambda_{\textnormal{speed}}^{2}>\lambda_{\textnormal{speed}}^{3}$, and $\phi_{t}^{i}$s, for $i\in\left\{1,2\right\}$, are handcrafted feature values encoding the obstacle distribution. Specially, $\phi_{t}^{i}$s are normalized linear combinations of the distance to the nearest obstacle, volume of the obstacle, and number of obstacles at current time .
In (10), $\phi_{t}^{1}$ is designed to shape the reward function smoother with a zero mean and $\phi_{t}^{2}$ is a naive estimate of the level of danger by human knowledge. Therefore, the first case in (10) serves as a dense smoothing of the collision penalty, which imposes the policy to behave conservatively when the observed environment is manually evaluated as dangerous. Such an approach prevents the policy from being single-mindedly greedy to gain instant reward by accelerating. Note that we do not necessarily carefully tune the hyperparameters in (10), but leave the task of further optimization to the objective reward in the second stage.
@@ V-D2 An objective reward for fine-tuning
In the second stage, we restore the original intent of the speed reward, which is only to encourage greater aggressiveness:
$r_{\textnormal{speed}}=\lambda_{\textnormal{speed}}^{3}\|\boldsymbol{v}_{t}\|.$ (11)
@@ VI-A System Setup
We choose a state-of-the-art model-based trajectory planner, EGO-planner [12], as the backbone. The speed constraint is imposed in the form of $\left\|\boldsymbol{v}\right\|\leq\bar{v}$. The local mapping algorithm is implemented based on [24]. The vehicle is equipped with a forward-facing camera that can measure depth with a FOV of 87 (horizontal) $\times$ 58 (vertical) and a valid/trusted depth range from $0.28$m to $5$m. The refreshing frequency of the depth data is 15 Hz. A PID controller running at 50 Hz is used for position tracking.
The outer-loop policy runs at 10 Hz. Conceptually, trajectory planning is required after each outer-loop policy output, but too frequent replanning increases the situations of triggering the emergency_stop state in practice, as well as increasing the computational overhead. Therefore, we design a criterion to determine whether replanning is imposed by the outer-loop policy output, and if this criterion is not met, the planner follows its original implementation of replanning rules [12]. In particular, if $v_{t}^{\dagger}-v_{t-1}^{\dagger}\in\left[-0.3,0.5\right]$m/s, we do not impose a replanning of the trajectory planner.
@@ VI-B Training Environment
We train the policy in a customized simulator, which parallelizes multiple (30 in our implementation) separate environments. Unlike general simulation environments, the simulator does not stop the clock while solving for the action, i.e., mapping and trajectory planning, in our simulation environment. Meanwhile, the sensory data is updated asynchronously with the action. These specialized designs are elaborated to simulate one of the most important factors limiting allowed aggressiveness, the perception latency, i.e., the time interval between perception and execution of an action. Even with these designs, however, we can only approximate perception latency because the computing platform at the deployment time is different from the one at the time of training. Another crucial factor that limits aggressiveness, the error of trajectory tracking due to limited performance of controller or possible dynamically infeasible trajectories, is also embedded in the environment by setting up the actual PID tracking controller for the vehicle, as described earlier.
@@ VI-C Training Implementation
Figure 4: Velocity distribution along the trajectory in different setups. The dark green curves are the reference trajectory. The colorful curves are the trajectory that the vehicle passes over, where the color represents the velocity. (a) An example result of the proposed approach. (b) An example result of the constant speed constraint ${v}^{\dagger}=2$m/s. (c) An example result of the intermediate model before fine-tuning.
Figure 5: Statistical results of different setups. Success rates are computed on 50 trials for each setup. The light blue dashed lines connect the statistics at different levels of constant speed constraint, indicating the inherent capability of the system. Average velocities are computed as the average of success trials in the 50 trials.
@@ VII-A Simulation Experiments
We benchmark three approaches in the simulation experiments: (i) the proposed ’learned policy + model-based planner backbone’, (ii) the backbone planner EGO-planner with constant speed constraint (i.e., constant ${v}^{\dagger}$), and (iii) the handcrafted cost function-based (non-learning) environmental adaptive planner EVA-planner [20], where the default parameters in the cost function are used, and all parameters that EGO-planner shares with it are set equal to those of EGO-planner.
@@
VII-A1 Comparison of the proposed framework with constant speed constraint baselines and ablation tests
The first two setups are evaluated in a challenging environment as shown in Fig. 4, where obstacles of different densities, sizes, and appearances are unevenly distributed. The evaluation environment is never seen when training. In Fig. 5, we report the statistical results of traversal velocity in success cases and the success rate which is defined as the proportion of trials that complete the entire track. A trial is judged to be ’terminal’ with the same criteria as during training. In Fig. 4, we visualize the trajectories and velocities at every moment of the proposed approach and the baselines. Since the trajectory generation backbone of EVA-planner cannot always effectively generate feasible trajectory in such a complex environment, it is not fair to include it in this statistical result, so we ignore EVA-planner in Fig. 5.
The statistics show that as ${v}^{\dagger}$ is set higher, the success rate decreases, creating a sloping downward performance curve in Fig. 5. Such performance curves capture the inherent performance of the integration of physical (e.g., sensors equipped) and algorithmic (e.g., the mapping, planning and control framework) systems. Although under the performance constraints of the system, by dynamically configuring the speed constriant, the vehicle combines the exploitation of agility and safety guarantees. It can be seen in Fig. 5 that the policy learned with both the reward stages achieves similar success rate as the lowest speed case where ${v}^{\dagger}=1.5$m/s, but also exhibits considerable traversal speeds to efficiently complete the track. In contrast, a relatively aggressive constant speed constraint ${v}^{\dagger}=2$m/s can frequently lead to emergency stops or collisions in areas with dense obstacles or local occlusion, as shown in Fig. 4b.
In Fig. 5, we also show the performance of the policies trained only with either the first stage (reward (10)) or the second stage (reward (11)). The results indicate that although the final policy can successfully improve the overall performance of the system, the policy trained with only the first stage does not exhibit such improvement. This result is not so surprising since the reward scheme of the first stage of training does not reflect the objective values and impose inaccurate human knowledge. On the other hand, the necessity of the first training stage is reflected by the inferiority of the performance rendered by the policy trained directly with sparse (but objective) rewards in Fig. 5.